An object with a mass of 0.5 kilogram is swung by a string in a horizontal circle of radius 1 meter at a speed of 2 meters per second. The magnitude of the force on the mass is


1
Expert's answer
2021-11-08T16:23:35-0500

"m=0.5Kg"

"v=2m\/s"

"r=1m"

"Force=mass\times\frac{(velocity)^2}{radius}"

"F=0.5\times\frac{(2)^2}{1}"

"F=2N."